OpenStudy (anonymous):

what is the acceleration experienced by a car that takes 10 seconds to reach 27 m/s from rest?

OpenStudy (anonymous):

Someone please help!!

OpenStudy (anonymous):

Vi= 0m/s Vf= 27m/s Delta T= 10S Use the formula Aave={(Vf-Vi)/2} (Delta T) Aave=[(27m/s-0m/s)/2](10s) Aave=135m/s2